# Locating Superspreaders in SI Cascades

### 05 Mar 2024 v1

@@ 8 Lower bounds for $\alpha\in(0,1/2)$
To prove the theorem, we will construct a tailored family of graphs ${\mathcal{H}}\subset{\mathcal{G}}$ for which we can tractably analyze the distribution of infection times, while also constituting a “hard" instance for estimating high-degree vertices. To construct ${\mathcal{H}}$, we first describe a particular deterministic graph $H$; see the definition below as well as Figure [todo] for a visualization.
@@
Fix a positive integer $N$. Let $H^{\prime}$ be any connected graph with $N$ vertices, maximum degree at most 3, and diameter at most $2\log N$ (e.g., a balanced binary tree). For each vertex in $u\in H^{\prime}$, we add a path of length $\log^{5}n$ starting at $u$; all added paths are disjoint. We call the resulting graph $H$. The set of endpoints of the paths added to $H^{\prime}$, which are also the set of leaves of $H$, is denoted by $L$.
@@
To construct the graphs of interest, we will introduce additional vertices to $H$ which connect to $L$ in a particular manner. See the formal definition below as well as the visualization in Figure [todo].
@@
  • If $G\in{\mathcal{H}}_{\emptyset}$ then $\mathsf{highdeg}(G)=\emptyset$, and if $G\in{\mathcal{H}}_{v}$ for $v\in L$ then $\mathsf{highdeg}(G)=v$.
@@
The first step in our analysis is to show that it is challenging to even detect whether the set of observed infection times are due to a graph in ${\mathcal{H}}_{\emptyset}$ or in ${\mathcal{H}}_{v}$ for some $v\in L$. To do so, we construct convenient distributions $\mu_{\emptyset}$ over elements of ${\mathcal{H}}_{\emptyset}$ and $\mu_{v}$ over elements of ${\mathcal{H}}_{v},v\in L$. We let $P_{\emptyset}$ and $P_{v},v\in L$ denote the induced distributions over the infection times. That is, for $x\in L\cup\{\emptyset\}$, $P_{x}$ is the distribution corresponding to $\mathcal{SI}(G,v_{0})^{\otimes K}$ where $G\sim\mu_{x}$. The precise description of the distributions $\{\mu_{x}\}_{x\in L\cup\{\emptyset\}}$ is somewhat complicated, so we defer the details to Section 8.1.
[play up the natural connections to the sparse mixture problem!]
@@
At a high level, we prove Lemma 8.5 by studying the $\chi^{2}$-divergence between $P_{\emptyset}$ and $P_{v}$ for $v\in L$, which provides an upper bound for the total variation distance and has a useful tensorization property which we leverage in our analysis. For full details, see Section 8.1.
With Lemma 8.5 in hand, we can now prove the theorem.
@@
Denote $\delta:=\max_{u,v\in L\cup\{\emptyset\}}\mathrm{TV}(Q_{u},Q_{v})$, and let $\hat{v}$ be any estimator. Then for any $u\in B\cup\{0\}$,
$P_{u}(\hat{v}=u)$ $=1-\sum_{v\in L\cup\{\emptyset\}:v\neq u}P_{u}(\hat{v}=v)$
$\leq 1-\sum_{v\in L\cup\{\emptyset\}:v\neq u}\left(P_{v}(\hat{v}=v)-\delta\right)$
$=1+\delta N-N\min_{v\in L\cup\{\emptyset\}}P_{v}(\hat{v}=v).$
As the upper bound holds for all $u\in L\cup\{\emptyset\}$, it follows that
$\min_{v\in L\cup\{\emptyset\}}P_{v}(\hat{v}=v)\leq 1+\delta N-N\min_{v\in L\cup\{\emptyset\}}P_{v}(\hat{v}=v).$
Rearranging terms shows that
$\min_{v\in L\cup\{\emptyset\}}P_{v}(\hat{v}=v)\leq\frac{1+\delta N}{1+N}\leq\frac{1}{N}+\delta=o(1),$ (8.1)
where the final inequality holds for $n$ sufficiently large in light of Lemma 8.5. Importantly, (8.1) implies the existence of $G\in{\mathcal{H}}$ for which $\operatorname{\mathbb{P}}_{G}(\hat{v}=\mathsf{highdeg}(G))\leq\epsilon$. As this conclusion holds for any estimator $\widehat{v}$, it holds that
$\sup_{\widehat{v}}\min_{G\in{\mathcal{G}}}\operatorname{\mathbb{P}}_{G}(\widehat{v}=v)=o(1),$
which is equivalent to the desired claim. ∎
8.1 Impossibility of detecting super-spreaders
8.1.1 Distributions of graphs in ${\mathcal{H}}$
Before proving the lemma, we must describe the distributions $\{\mu_{x}\}_{x\in L\cup\{\emptyset\}}$ over graphs in ${\mathcal{H}}$. To this end, it is useful to first define corresponding “auxiliary" distributions $\{\mu_{x}^{\prime}\}_{x\in L\cup\{\emptyset\}}$. We say that $G\sim\mu_{\emptyset}^{\prime}$ if each vertex in $[N]$ chooses a uniform random element of $L$ with which they form an edge. On the other hand, we say that $G\sim\mu_{v}^{\prime}$ if, for each $u\in[N]$, $u$ connects to a uniform random element of $L$ with probability $1-D/N$, otherwise $u$ connects to $v$ with probability $D/N$. Finally, for $x\in L\cup\{\emptyset\}$, the measure $\mu_{x}$ is equal to $\mu_{x}^{\prime}$ conditioned on the sampled graph being an element of ${\mathcal{H}}_{x}$.
@@ 8.1.1 Distributions of graphs in ${\mathcal{H}}$
8.1.2 Proof of Lemma 8.5
The advantage of using the auxiliary distributions is that they are product distributions conditioned on the infection times in $H$, which makes it easier to analyze the infection times in $[N]$. To see this concretely, let
@@
The total variation distance and the $\chi^{2}$ divergence can be related through the inequality [polyanskiy_wu_information_theory, Proposition 7.15]
@@
For our setting, the $\chi^{2}$ divergence has two advantageous properties. First, it tensorizes; that is, for any positive integer $m$, it holds that [polyanskiy_wu_information_theory, Chapter 7.12]
$\chi^{2}\left(\left.P^{\otimes m}\right\|Q^{\otimes m}\right)\leq\left(1+\chi^{2}(P\|Q)\right)^{m}-1.$
@@
We are now ready to complete the proof of the main result.
@@
In particular, for any $\epsilon>0$ the condition (8.10) is satisfied if
@@
Above, the inequality $(a)$ follows from the property that conditioning increases the total variation [polyanskiy_wu_information_theory, Theorem 7.5], and inequality $(b)$ holds since the total variation is always upper bounded by 1. Finally, $\mathrm{TV}(P_{\emptyset}^{\prime},P_{v}^{\prime})=o(1)$ implies that $\mathrm{TV}(P_{\emptyset},P_{v})=o(1)$ as well in light of (8.2). ∎
8.2 Analysis of the likelihood ratio: Proof of Lemma 8.7
In this section, we study the likelihood ratio
@@
To formalize these ideas, we begin by defining the following “nice" event.
@@
We let ${\mathcal{E}}_{v,1},{\mathcal{E}}_{v,2}$ and ${\mathcal{E}}_{v,3}$ denote the events corresponding to the first, second and third conditions of ${\mathcal{E}}_{v}$ in Definition 8.8. We start by showing that ${\mathcal{E}}_{v,1}$ holds with probability $1-o(1)$. By our assumption that the diameter of $H^{\prime}$ is at most $2\log N$ (see Definition 8.2), we have that $T_{i}(z)$ is stochastically upper bounded by a $\mathrm{Gamma}(2\log N,1)$ random variable. Hence
@@
$\operatorname{\mathbb{P}}({\mathcal{E}}_{v,3}^{c})=\operatorname{\mathbb{P}}\left(\sum_{i=1}^{K}(X_{i}(v)-R)^{2}>2K\right)\leq\frac{1}{4K^{2}R^{2}}\operatorname{Var}\left(\sum_{i=1}^{K}(X_{i}(v)-R)^{2}\right)\leq\frac{2}{K}=o(1).$ (8.18)
@@
Throughout this proof, we shall condition on the infection times in $H^{\prime}$, given by the collection $\{\mathbf{T}(z)\}_{z\in H^{\prime}}$, as well as $v$, given by $\mathbf{T}(v)$. We will also assume that the “nice" event ${\mathcal{E}}_{v}$ holds, which is measurable with respect to $\{\mathbf{T}(z)\}_{z\in H^{\prime}},\mathbf{T}(v)$. Finally, as a shorthand, we will denote $R:=\log^{5}N$.
Recall from Fact 8.3 that $\{\mathbf{T}(u)\}_{u\in L}$ is a collection of conditionally independent variables with respect to $\{\mathbf{T}(z)\}_{z\in H^{\prime}}$. As a result, if we condition on both $\{\mathbf{T}(z)\}_{z\in H^{\prime}}$ and $\mathbf{T}(v)$, then the terms
@@
with probability at least $1-2\exp(-2N^{1/3})$. We proceed by studying the conditional expectation in (8.20). To this end, for any $w\in L$, we denote $w^{\prime}$ to be the last vertex in $H^{\prime}$ on the unique path starting from $v_{0}$ and ending at $w$. We have that
@@
$\operatorname{\mathbb{E}}\left[\left.\frac{1}{|B|}\sum_{w\in B}\prod_{i=1}^{K}e^{T_{i}(w)-T_{i}(v)}\mathbf{1}(T_{i}(v)\geq T_{i}(w))\right|\{\mathbf{T}(z)\}_{z\in H^{\prime}},\mathbf{T}(v)\right]\\ \geq\left(\frac{e^{-16}}{3\sqrt{R}}\right)^{K}\exp\left\{-\frac{4}{R}\sum_{i=1}^{K}(X_{i}(v)-R)^{2}\right\}\stackrel{{\scriptstyle(f)}}{{\geq}}\left(\frac{e^{-30}}{\sqrt{R}}\right)^{K},$ (8.23)
@@
$\frac{1}{|B|}\sum_{w\in B}\prod_{i=1}^{K}e^{T_{i}(w)-T_{i}(v)}\mathbf{1}(T_{i}(v)\geq T_{i}(w))\geq\left(\frac{e^{-30}}{\sqrt{R}}\right)^{K}-N^{-1/3}\geq R^{-K}.$
@@
$\operatorname{\mathbb{P}}\left(\left.|I[t,t+\epsilon]|\geq 2\right|{\mathcal{F}}_{t}\right)=o(\epsilon).$
@@
Recall that for a subset of vertices $S\subset V$, we denote $\operatorname{\mathsf{cut}}(S):=\{(u,w)\in E(G):u\in S,v\notin S\}$. We also define $\tau_{1}$ and $\tau_{2}$ to be the first and second times after $t$ when a new vertex is added to the cascade, respectively. Conditionally on ${\mathcal{F}}_{t}$, the memoryless property of Exponential distributions implies that $\tau_{1}-t$ is equal in distribution to $\mathrm{Exp}(\lambda|\operatorname{\mathsf{cut}}(I[0,t])|)$, which stochastically dominates a $\mathrm{Exp}(\lambda n^{2})$ random variable (here, we have used that the number of edges in $G$ is at most $n^{2}$). By the memoryless property of Exponential distributions, $\tau_{2}-\tau_{1}$ is independent of $\tau_{1}-t$. Through similar argument as above, it holds that $\tau_{2}-\tau_{1}$ stochastically dominates a $\mathrm{Exp}(\lambda n^{2})$ random variable as well.
Putting everything together, the probability that at least two vertices join the cascade in the time interval $[t,t+\epsilon]$ is
@@
We start by defining some relevant notation. Condition on ${\mathcal{F}}_{t}$, and let $v\in V\setminus I[0,t]$. Recalling that $\{F_{t}(e)\}_{e\in E(G)}$ is a collection of i.i.d. $\mathrm{Exp}(\lambda)$ random variables, we define
$W$ $:=\min\{F_{t}(w,v):w\in I[0,t],(w,v)\in E(G)\}.$
By basic properties of Exponential distributions, we have that $W\sim\mathrm{Exp}(\lambda|{\mathcal{N}}(v)\cap I[0,t]|)$. Additionally define the event ${\mathcal{A}}$ to hold if and only if at most one vertex joins the cascade in the interval $(t,t+\epsilon]$.
We claim that if ${\mathcal{A}}$ holds, then $v\in I[0,t+\epsilon]\iff W\leq\epsilon$. Indeed, if $W\leq\epsilon$, then from the description of $T(v)$ in the proposition statement, we have that $T(v)\leq t+\epsilon$, which implies that $v\in I[t,t+\epsilon]$. On the other hand, if $W>\epsilon$, then we consider two cases: (1) no vertex becomes infected in $[t,t+\epsilon]$ or (2) one vertex becomes infected in $[t,t+\epsilon]$. Note that no other case is possible if ${\mathcal{A}}$ holds. In the first case, it holds for all $e\in\operatorname{\mathsf{cut}}(I[0,t])$ that $F_{t}(e)>\epsilon$, hence the weight of any path starting in $I[0,t]$ and ending at $v$ is at least $\epsilon$, implying that $T(v)>t+\epsilon$. In the second case, if $v\in I[t,t+\epsilon]$ despite $W>\epsilon$, there must exist a path $P\in{\mathcal{P}}(v,t)$ of length at least 2 such that $\mathsf{weight}(P,t)\leq\epsilon$. However, this implies the existence of another vertex $u$ on the path which is not an element of $I[0,t]$. Since the portion of the path halting at $u$ has a smaller weight than $P$, it follows that $T(u)\in I[t,t+\epsilon]$. However, it is not possible for both $u$ and $v$ to be elements of $I[t,t+\epsilon]$ if ${\mathcal{A}}$ holds, hence we must have that $T(v)>t+\epsilon$.
@@
($\operatorname{\mathbb{P}}\left(\{v\in I[0,t+\epsilon]\}\cap{\mathcal{A}}|{\mathcal{F}}_{t}\right)$ =($=\operatorname{\mathbb{P}}\left(\{W\leq\epsilon\}\cap{\mathcal{A}}|{\mathcal{F}}_{t}\right)$
$=\operatorname{\mathbb{P}}(W\leq\epsilon|{\mathcal{F}}_{t})-\operatorname{\mathbb{P}}\left(\{W\leq\epsilon\}\cap{\mathcal{A}}^{c}|{\mathcal{F}}_{t}\right)$
$=\epsilon\lambda|{\mathcal{N}}(v)\cap I[0,t]|+o(\epsilon).$
Finally, since $\operatorname{\mathbb{P}}({\mathcal{A}}|{\mathcal{F}}_{t})=1-o(\epsilon)$, it holds that
$\operatorname{\mathbb{P}}\left(v\in I[0,t+\epsilon]|{\mathcal{F}}_{t}\right)=\epsilon\lambda|{\mathcal{N}}(v)\cap I[0,t]|+o(\epsilon),$

### 06 Mar 2024 v2

@@ 3 Preliminaries: Cascade dynamics and infection times
To prove the theorem, we will construct a tailored family of graphs ${\mathcal{H}}\subset{\mathcal{G}}$ for which we can tractably analyze the distribution of infection times, while also constituting a “hard” instance for estimating high-degree vertices. To construct ${\mathcal{H}}$, we first describe a particular deterministic graph $H$.
@@
Fix a positive integer $N$. Let $H^{\prime}$ be a connected tree with $N$ vertices, maximum degree at most 3, and diameter at most $2\log N$ (e.g., a balanced binary tree). For each vertex in $u\in H^{\prime}$, we add a path of length $\log^{5}n$ starting at $u$; all added paths are disjoint. We call the resulting graph $H$. The set of endpoints of the paths added to $H^{\prime}$, which are also the set of leaves of $H$, is denoted by $L$.
@@
To construct the graphs of interest, we will introduce additional vertices to $H$ which connect to $L$ in a particular manner.
@@
  • If $G\in{\mathcal{H}}_{\emptyset}$ then $\mathsf{highdeg}(G)=\emptyset$, and if $G\in{\mathcal{H}}_{v}$ for $v\in L$ then $\mathsf{highdeg}(G)=\{v\}$.
@@
[…]
@@
At a high level, we prove Lemma 6.4 by studying the $\chi^{2}$-divergence between $P_{\emptyset}$ and $P_{v}$ for $v\in L$, which provides an upper bound for the total variation distance and has a useful tensorization property which we leverage in our analysis. Before proving the lemma, we show how it can be used to prove the theorem.
@@
Define the measure $\mu$ over the ensemble ${\mathcal{H}}$ via $\mu:=\frac{1}{N+1}\sum_{x\in L\cup\{\emptyset\}}\mu_{x}$, and similarly define the induced measure over vertex infection times $P:=\frac{1}{N+1}\sum_{x\in L\cup\{\emptyset\}}P_{x}$. Let us also define $\delta:=\max_{x,y\in L\cup\{\emptyset\}}\mathrm{TV}(P_{x},P_{y})$ and note that $\delta=o(1)$ by Lemma 6.4 and the triangle inequality. For any estimator $\mathrm{HD}$, it holds that
$P(\mathrm{HD}=\mathsf{highdeg}(G))$ $=\frac{1}{N+1}\sum_{x\in L\cup\{\emptyset\}}P_{x}(\mathrm{HD}=\{x\})$
=$\stackrel{{\scriptstyle(a)}}{{=}}1-\frac{1}{N+1}\sum_{x,y\in L\cup\{\emptyset\}:x\neq y}P_{x}(\mathrm{HD}=\{y\})$
$\stackrel{{\scriptstyle(b)}}{{\leq}}1-\frac{1}{N+1}\sum_{x,y\in L\cup\{\emptyset\}:x\neq y}\left(P_{y}(\mathrm{HD}=\{y\})-\delta\right)$
=$\stackrel{{\scriptstyle(c)}}{{=}}1+N\delta-NP(\mathrm{HD}=\mathsf{highdeg}(G)),$
where $(a)$ is due to the law of total probability, $(b)$ uses the definition of the total variation distance, and $(c)$ follows as each term $P_{y}(\mathrm{HD}=\{y\})$ is counted $N$ times in the summation. Rearranging terms shows that
$P(\mathrm{HD}=\mathsf{highdeg}(G))\leq\frac{1+N\delta}{1+N}=o(1).$
The remainder of this section is devoted to the proof of Lemma 6.4.
6.1 Distributions of graphs in ${\mathcal{H}}$
Before proving the lemma, we must describe the distributions $\{\mu_{x}\}_{x\in L\cup\{\emptyset\}}$ over graphs in ${\mathcal{H}}$. To this end, it is useful to first define corresponding “auxiliary” distributions $\{\mu_{x}^{\prime}\}_{x\in L\cup\{\emptyset\}}$. We say that $G\sim\mu_{\emptyset}^{\prime}$ if each vertex in $[N]$ chooses a uniform random element of $L$ with which they form an edge. On the other hand, we say that $G\sim\mu_{v}^{\prime}$ if, for each $u\in[N]$, $u$ connects to a uniform random element of $L$ with probability $1-D/N$, otherwise $u$ connects to $v$ with probability $D/N$. Finally, for $x\in L\cup\{\emptyset\}$, the measure $\mu_{x}$ is equal to $\mu_{x}^{\prime}$ conditioned on the sampled graph being an element of ${\mathcal{H}}_{x}$.
@@ 6.1 Distributions of graphs in ${\mathcal{H}}$
6.2 Impossibility of detection: Proof of Lemma 6.4
The advantage of using the auxiliary distributions is that they are product distributions conditioned on the infection times in $H$, which makes it easier to analyze the infection times in $[N]$. To see this concretely, let us assume without loss of generality that $\lambda=1$, and define
@@
The total variation distance and the $\chi^{2}$ divergence can be related through the inequality [41, Proposition 7.15]
@@
For our setting, the $\chi^{2}$ divergence has two advantageous properties. First, it tensorizes; that is, for any positive integer $k$, it holds that [41, Chapter 7.12]
$\chi^{2}\left(\left.P^{\otimes k}\right\|Q^{\otimes k}\right)=\left(1+\chi^{2}(P\|Q)\right)^{k}-1.$
@@
Before proving Lemma 6.6, we show how it can be used to prove the main result.
@@
Since $D/N=n^{\alpha-1+o(1)}$, for any $\epsilon>0$ the condition (6.9) is satisfied if
@@
Above, the inequality $(a)$ follows from the property that conditioning increases the total variation [41, Theorem 7.5], and inequality $(b)$ holds since the total variation is always upper bounded by 1. Finally, $\mathrm{TV}(P_{\emptyset}^{\prime},P_{v}^{\prime})=o(1)$ implies that $\mathrm{TV}(P_{\emptyset},P_{v})=o(1)$ as well in light of (6.1). ∎
In the remainder of this subsection, we study the likelihood ratio
@@
To formalize these ideas, we begin by defining the following “nice” event.
@@
We let ${\mathcal{E}}_{v,1},{\mathcal{E}}_{v,2}$ and ${\mathcal{E}}_{v,3}$ denote the events corresponding to the first, second and third conditions of ${\mathcal{E}}_{v}$ in Definition 6.7. We start by showing that ${\mathcal{E}}_{v,1}$ holds with probability $1-o(1)$. Since $H^{\prime}$ is a tree and the diameter of $H^{\prime}$ is at most $2\log N$ (see Definition 6.1), we have that $T_{i}(z)$ is stochastically upper bounded by a $\mathrm{Gamma}(2\log N,1)$ random variable. Hence
@@
$\operatorname{\mathbb{P}}({\mathcal{E}}_{v,3}^{c})=\operatorname{\mathbb{P}}\left(\sum_{i=1}^{K}(X_{i}(v)-R)^{2}>2KR\right)\leq\frac{1}{K^{2}R^{2}}\operatorname{Var}\left(\sum_{i=1}^{K}(X_{i}(v)-R)^{2}\right)\leq\frac{8}{K}=o(1).$ (6.17)
@@
Throughout this proof, we shall condition on the infection times in $H^{\prime}$, given by the collection $\{\mathbf{T}(z)\}_{z\in H^{\prime}}$, as well as $v$, given by $\mathbf{T}(v)$. We will also assume that the “nice” event ${\mathcal{E}}_{v}$ holds, which is measurable with respect to $\{\mathbf{T}(z)\}_{z\in H^{\prime}},\mathbf{T}(v)$. Finally, as a shorthand, we will denote $R:=\log^{5}N$.
Recall from Fact 6.2 that $\{\mathbf{T}(u)\}_{u\in L}$ is a collection of conditionally independent variables with respect to $\{\mathbf{T}(z)\}_{z\in H^{\prime}}$. As a result, if we condition on both $\{\mathbf{T}(z)\}_{z\in H^{\prime}}$ and $\mathbf{T}(v)$, then the terms
@@
with probability at least $1-2\exp(-2N^{1/3})$. We proceed by studying the conditional expectation in (6.19). To this end, for any $w\in L$, we denote $w^{\prime}$ to be the last vertex in $H^{\prime}$ on the unique path starting from $v_{0}$ and ending at $w$; such a vertex exists since $H$ is connected. We have that
@@
$\operatorname{\mathbb{E}}\left[\left.\frac{1}{N}\sum_{w\in L}\prod_{i=1}^{K}e^{T_{i}(w)-T_{i}(v)}\mathbf{1}(T_{i}(v)\geq T_{i}(w))\right|\{\mathbf{T}(z)\}_{z\in H^{\prime}},\mathbf{T}(v)\right]\\ \geq\left(\frac{e^{-16}}{3\sqrt{R}}\right)^{K}\exp\left\{-\frac{4}{R}\sum_{i=1}^{K}(X_{i}(v)-R)^{2}\right\}\stackrel{{\scriptstyle(f)}}{{\geq}}\left(\frac{e^{-30}}{\sqrt{R}}\right)^{K},$ (6.22)
@@
$\frac{1}{N}\sum_{w\in L}\prod_{i=1}^{K}e^{T_{i}(w)-T_{i}(v)}\mathbf{1}(T_{i}(v)\geq T_{i}(w))\geq\left(\frac{e^{-30}}{\sqrt{R}}\right)^{K}-N^{-1/3}\geq R^{-K}.$
@@
$\operatorname{\mathbb{P}}\left(\left.|{\mathcal{I}}[t,t+\epsilon]|\geq 2\right|{\mathcal{F}}_{t}\right)=o(\epsilon).$
@@
Define $\tau_{1}$ and $\tau_{2}$ to be the first and second times after $t$ when a new vertex is added to the cascade, respectively. Conditionally on ${\mathcal{F}}_{t}$, the memoryless property of Exponential distributions implies that $\tau_{1}-t$ is equal in distribution to $\mathrm{Exp}(\lambda|\operatorname{\mathsf{cut}}({\mathcal{I}}[0,t])|)$, which stochastically dominates a $\mathrm{Exp}(\lambda n^{2})$ random variable (here, we have used that the number of edges in $G$ is at most $n^{2}$). By the memoryless property of Exponential distributions, $\tau_{2}-\tau_{1}$ is independent of $\tau_{1}-t$. Through similar argument as above, it holds that $\tau_{2}-\tau_{1}$ stochastically dominates a $\mathrm{Exp}(\lambda n^{2})$ random variable as well. Consequently, the probability that at least two vertices join the cascade in the time interval $[t,t+\epsilon]$ is
@@
We start by defining some relevant notation. Condition on ${\mathcal{F}}_{t}$, and let $v\in V\setminus{\mathcal{I}}[0,t]$. Recalling that $\{F(e)\}_{e\in E}$ is a collection of i.i.d. $\mathrm{Exp}(\lambda)$ random variables, we define
$W$ $:=\min\{F(w,v):w\in{\mathcal{I}}[0,t],(w,v)\in E\}.$
By basic properties of Exponential distributions, we have that $W\sim\mathrm{Exp}(\lambda|{\mathcal{N}}(v)\cap{\mathcal{I}}[0,t]|)$. Additionally define the event ${\mathcal{E}}$ to hold if and only if at most one vertex joins the cascade in the interval $(t,t+\epsilon]$.
We claim that if ${\mathcal{E}}$ holds, then $v\in{\mathcal{I}}[0,t+\epsilon]\iff W\leq\epsilon$. Indeed, if $W\leq\epsilon$, then from the description of $T(v)$ in the proposition statement, we have that $T(v)\leq t+\epsilon$, which implies that $v\in{\mathcal{I}}[t,t+\epsilon]$. On the other hand, if $W>\epsilon$, then we consider two cases: (1) no vertex becomes infected in $[t,t+\epsilon]$ or (2) one vertex becomes infected in $[t,t+\epsilon]$. Note that no other case is possible if ${\mathcal{E}}$ holds. In the first case, it holds for all $e\in\operatorname{\mathsf{cut}}({\mathcal{I}}[0,t])$ that $F(e)>\epsilon$, hence the weight of any path starting in ${\mathcal{I}}[0,t]$ and ending at $v$ is at least $\epsilon$, implying that $T(v)>t+\epsilon$. In the second case, if $v\in{\mathcal{I}}[t,t+\epsilon]$ despite $W>\epsilon$, there must exist a path $P\in{\mathcal{P}}(v,t)$ of length at least 2 such that $\mathsf{weight}(P,t)\leq\epsilon$. However, this implies the existence of another vertex $u$ on the path which is not an element of $I[0,t]$. Since the portion of the path halting at $u$ has a smaller weight than $P$, it follows that $T(u)\in{\mathcal{I}}[t,t+\epsilon]$. However, it is not possible for both $u$ and $v$ to be elements of ${\mathcal{I}}[t,t+\epsilon]$ if ${\mathcal{E}}$ holds, hence we must have that $T(v)>t+\epsilon$.
@@
($\operatorname{\mathbb{P}}\left(\{v\in{\mathcal{I}}[0,t+\epsilon]\}\cap{\mathcal{E}}|{\mathcal{F}}_{t}\right)$ =($=\operatorname{\mathbb{P}}\left(\{W\leq\epsilon\}\cap{\mathcal{E}}|{\mathcal{F}}_{t}\right)$
$=\operatorname{\mathbb{P}}(W\leq\epsilon|{\mathcal{F}}_{t})-\operatorname{\mathbb{P}}\left(\{W\leq\epsilon\}\cap{\mathcal{E}}^{c}|{\mathcal{F}}_{t}\right)$
$=\epsilon\lambda|{\mathcal{N}}(v)\cap{\mathcal{I}}[0,t]|+o(\epsilon).$
Finally, since $\operatorname{\mathbb{P}}({\mathcal{E}}|{\mathcal{F}}_{t})=1-o(\epsilon)$ by Lemma 7.1, it holds that
$\operatorname{\mathbb{P}}\left(v\in{\mathcal{I}}[0,t+\epsilon]|{\mathcal{F}}_{t}\right)=\epsilon\lambda|{\mathcal{N}}(v)\cap{\mathcal{I}}[0,t]|+o(\epsilon),$